OpenStudy (anonymous):

A cone has a volume of 24 cubic inches. What is the volume of a cylinder that the cone fits exactly inside of?

OpenStudy (campbell_st):

the volume of a cone is \[V = \frac{1}{3} \pi r^2h\] volume of a cylinder is \[V = \pi r^2h\] so if they both have the same height and radius.... then the formula says the cone is 1/3 of the cylinder... so multiply the cone by 3 to get the volume of the cylinder